### v1

# Enhancing Graph-based Recommendations with Majority-Voting LLM-Rerank Augmentation

Abstract
Recommendation systems often suffer from data sparsity caused by limited user-item interactions, which degrade their performance and amplify popularity bias in real-world scenarios. This paper proposes a novel data augmentation framework that leverages Large Language Models (LLMs) and item textual descriptions to enrich interaction data. By few-shot prompting LLMs multiple times to rerank items and aggregating the results via majority voting, we generate high-confidence synthetic user-item interactions, supported by theoretical guarantees based on the concentration of measure. To effectively leverage the augmented data in the context of a graph recommendation system, we integrate it into a graph contrastive learning framework to mitigate distributional shift and alleviate popularity bias. Extensive experiments show that our method improves accuracy and reduces popularity bias, outperforming strong baselines.
1 Introduction
Recommendation systems (RS), particularly graph-based recommendation models, often suffer from data sparsity, which significantly degrades performance and amplifies popularity bias (Idrissi and Zellou 2020). Traditional data augmentation techniques such as random walk sampling, node (edge) dropout (Wu et al. 2021), and subgraph sampling (Fan et al. 2023a; Li et al. 2023) generate alternative views of the interaction graph for use in self-supervised learning. These methods are simple and can improve accuracy in low-interaction regimes, but they rely solely on interaction data and overlook the potential of other modalities, such as item textual descriptions. Recent advances leverage the reasoning capabilities of Large Language Models (LLMs) and the rich semantic information in item descriptions to enhance user and item profiling for data augmentation (Wu et al. 2024b).
While LLM-based methods show promise in enriching recommendation datasets using external knowledge, they suffer from two main limitations. First, the augmentation results generated by prompting LLMs are often inconsistent across different inference runs, leading to unstable performance. (Hou et al. 2024b; Wei et al. 2024). Specifically, using a commonly adopted prompt template from prior research to augment interaction data, the results on the Netflix dataset sampled with 100010001000 users to reduce augmentation cost show that LLMRec (Hou et al. 2024b) and Llama4Rec (Luo et al. 2024) yield inconsistent NDCG@10 scores. This metric evaluates how well the ranking of the top 10 items reflects user preferences, as illustrated in Figure 1.
Figure 1: NDCG@10 changes over LLM augmentation runs
Figure 2: NDCG@10 of LLM-based reranking over multiple inference runs. Red dashed line: LightGCN baseline; green solid line: ensemble result of LLM runs on 2 datasets.
This is because the augmented items vary across different runs, leading to unstable outcomes. Second, embeddings derived from the textual content of LLM-embedded models tend to be out-of-distribution compared to collaborative embeddings, making it difficult to align them in a model-agnostic framework across various RS (Ren et al. 2024; Qiao et al. 2024).
To address these issues, we propose a reformulated prompting strategy that casts the augmentation task as an item reranking problem, enhanced with a few-shot prompting design. This format enables the LLM to better leverage its parametric knowledge when ranking candidate items according to user preferences, encouraging the model to assign high positive scores to items that are more likely to match user interests (Wu et al. 2024a; Hou et al. 2024c). As shown in Figure 2, although individual LLM reranking runs yield varying NDCG@10 scores, they outperform the traditional graph-based method LightGCN (He et al. 2020) in most cases. Aggregating multiple reranking outputs via majority voting further improves and stabilizes performance, demonstrating the LLM’s capacity to adapt to user preferences. To quantify this effect, we vary the number of voting runs and conduct 30 repeated inferences for each setting.
Figure 3: Distribution of NDCG@10 over different number of ensemble
The results, visualized in Figure 3, show that increased voting numbers improves mean NDCG@10 and reduces variance, indicating greater robustness.
Motivated by this, we propose a theoretically grounded data augmentation strategy based on majority-vote LLM reranking, supported by concentration of measure theory to bound augmentation accuracy. Instead of incorporating LLM-generated textual embeddings, we directly leverage the augmented data via a contrastive learning (CL) objective applied between original and augmented graphs. This enables representation alignment while mitigating distributional shift and popularity bias, without modifying the base architecture. We summarize our main contributions as follows: (1), we propose VoteGCL, a model-agnostic, end-to-end framework that employs few-shot prompted LLMs for in-context reranking. By performing majority-vote reranking over candidate items, VoteGCL identifies high-confidence user-item interactions for data augmentation. It further integrates graph contrastive learning (CL) between original and augmented views to exploit the augmented data and address distributional shift; (2) we provide a theoretical analysis based on the concentration of measure, demonstrating that majority-vote reranking effectively extracts the LLM’s most reliable knowledge, thereby enhancing the quality of augmented interactions; and (3) extensive experiments show that VoteGCL consistently outperforms state-of-the-art graph-based, multi-modal, and LLM-enhanced recommender models across multiple benchmarks, improving recommendation accuracy while mitigating popularity bias.
2 Related work
Graph-Augmented Recommendation. Graph-based RS represent user-item interactions as bipartite graphs, where nodes are users and items, and edges denote interactions. Early models used random walks to uncover latent relationships via transition probabilities (Nikolakopoulos and Karypis 2019). The introduction of Graph Convolutional Networks (GCNs) enabled embedding propagation to better capture collaborative signals (He et al. 2020; Wang et al. 2019). Various graph augmentation techniques have been developed to address data sparsity and over-smoothing issues in graph neural networks. Traditional graph augmentation methods (Ding et al. 2022) can be categorized into three approaches: (1) Edge/Node Dropout, which removes redundant connections to improve robustness (Wu et al. 2021; Zhou et al. 2023); (2) Graph Diffusion, which enriches connectivity by adding new edges based on structural properties (Fan et al. 2023b; Yang et al. 2021); and (3) Subgraph Sampling, which generates localized views for diverse training samples (Zhang et al. 2024; Xie et al. 2022). While effective, these methods focus solely on graph structure and overlook valuable side information such as item content, motivating more holistic augmentation strategies.
LLMs-Enhanced Graph Recommendation. Recent advances explore integrating large language models (LLMs) into RS, typically along three lines: (1) Knowledge Enhancement, which uses LLMs’ reasoning capabilities to generate comprehensive user profiles and item descriptions by summarizing user behavior patterns and extracting rich semantic features from item content (Liu et al. 2024; Xi et al. 2024; Wang et al. 2024); (2) Model Enhancement, where LLM-generated textual embeddings of item descriptions and user profiles are integrated with collaborative signals through attention mechanisms or multi-modal alignment (Ren et al. 2024; Qiao et al. 2024); and (3) Interaction Enhancement employs LLMs to augment the user-item interaction space by generating synthetic interactions or directly scoring candidate items. This includes predicting missing ratings, generating counterfactual interactions to address data sparsity (Wei et al. 2024; Song, Chao, and Liu 2024). Despite their promise, LLM-based methods face challenges such as representation misalignment and inconsistent inference due to the stochastic nature of generation.
3 Preliminaries
Figure 4: Overview of the VoteGCL. The framework starts with a retrieval model that generates candidate items and identifies similar users for low-degree users using collaborative signals. An LLM reranks these candidates, and majority voting selects top-K high-confidence interactions for augmentation. The augmented data is then integrated into a graph contrastive learning framework, where two contrastive views are created via separate graph convolutions on a shared embedding matrix.
Problem Formulation. In graph recommendation, user-item interactions are modeled as a bipartite graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\mathcal{U}\cup\mathcal{I}$ includes user nodes $\mathcal{U}$ and item nodes $\mathcal{I}$, and $\mathcal{E}$ denotes observed interactions. Each user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$ is associated with a learnable embedding $\mathbf{e}_{u},\mathbf{e}_{i}\in\mathbb{R}^{d}$, initialized in the embedding table $\mathbf{E}^{(0)}\in\mathbb{R}^{|\mathcal{V}|\times d}$ and updated during training, where $d$ is the embedding size. The training process is performed on mini-batches $\mathcal{B}$ of user-item interactions sampled from $\mathcal{E}$.
Vanilla Graph-based Retrieval. We adopt LightGCN (He et al. 2020), a simplified yet effective GCN, as the primary retrieval model to propagate information over the user-item graph $\mathcal{G}$ by updating node representations through neighborhood aggregation and combination. At the $l$-th iteration, the node representations are updated as: $\mathbf{E}^{(l)}=H(\mathbf{E}^{(l-1)},\mathcal{G}),$ where $\mathbf{E}^{(l)}$ denotes the node representations at iteration $l$ which corresponds to the $l$-th layer of graph convolution, and $H$ is the neighborhood aggregation function. The embedding propagation process is mathematically formulated as a weighted combination of representations from different iterations. Specifically, the final node embeddings are computed as:
$\mathbf{E}=\frac{1}{1+L}\left(\mathbf{E}^{(0)}+\tilde{\mathbf{A}}\mathbf{E}^{(0)}+\ldots+\tilde{\mathbf{A}}^{L}\mathbf{E}^{(0)}\right),$ (1)
where $\mathbf{E}$ denotes the final representations used for downstream prediction tasks, $L$ corresponds to the number of propagation layers, and $\tilde{\mathbf{A}}\in\mathbf{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is the normalized undirected adjacency matrix without self-connections. This iterative process allows the model to capture higher-order connectivity patterns in the graph. A common approach is to train the graph-based recommendation model using the Bayesian Personalized Ranking (BPR) loss. The model learns to assign higher scores to observed interactions than to unobserved ones. The predicted score for user $u$ and item $i$ is computed via inner product: $\hat{y}_{ui}=\mathbf{e}_{u}^{\top}\mathbf{e}_{i}.$ The BPR loss is defined as:
$\mathcal{L}_{\text{BPR}}=\sum_{(u,i,j)\in\mathcal{O}}-\log\sigma\left(\hat{y}_{ui}-\hat{y}_{uj}\right),$ (2)
where $O=\left\{(u,i,j)\mid(u,i)\in\mathcal{E},(u,j)\in\mathcal{R}^{-}\right\}$ denotes the pairwise training data, $\mathcal{R}^{-}$is the unobserved interactions and $\sigma(\cdot)$ is the sigmoid function. During training, the user embedding $\mathbf{e}_{u}$ is optimized to be close to the embeddings $\mathbf{e}_{i}$ of preferred items. In inference, recommendations are made by selecting items whose embeddings are nearest to $\mathbf{e}_{u}$.
4 Methodology
We propose framework VoteGCL, as illustrated in Figure 4, we use few-shot prompting LLM for inference via majority-voted reranking to construct the augmented dataset as detailed in Section 4.1. To effectively leverage the augmented data for graph recommendation, we employ the CL paradigm as detailed in Section 4.2.
4.1 Data Augmentation via Majority LLM Reranking
Augmenting artificial interactions for high-degree users can introduce unnecessary cost and noisy interactions. Therefore, we apply the augmentation only to low-degree users, defined as those whose number of interactions falls below the 25th percentile ($Q_{\alpha}=0.25$) of the user interaction distribution. This strategy ensures that edges are augmented only for users with insufficient interaction data. An ablation study analyzing the effect of this threshold is provided in Section 5.3.
For each user $u$ in the selected set, we aim to select $p$ new interactions that do not appear in the original dataset. On the other hand, while LLMs have demonstrated greater effectiveness, they are limited by context length and thus cannot consider all potential items and related users within a single prompt. Hence, we first propose applying a conventional retrieval model to obtain an ordered shortlist of $K$ non-interacted candidate items, denoted as $\textbf{I}_{\text{can}}=\{i_{k}\}_{k=1}^{K}$, along with a set of similar users based on collaborative embeddings generated by the retrieval model, inspired by Wu et al.; Wei et al. Based on this information, we employ an LLM to rerank the sequence $\textbf{I}_{\text{can}}$, producing a sorted list denoted as $\textbf{I}_{\text{sorted}}$. To perform the reranking, we design a generalizable few-shot prompting template consisting of four components: first, a task-specific instruction; second, the user’s watch history enriched with flattened metadata; third, a structured output format; and fourth, one or more few-shot examples drawn from similar users. Each example includes the user’s history along with a ground-truth reranked list sorted by rating. This prompt is dataset-agnostic and can be easily adapted to new domains by reformatting the metadata. An example prompt for the movie dataset is provided in Figure 4, with further details available in Appendix LABEL:prompt.
To improve the consistency of reranking outcomes and better leverage the LLM’s reasoning capabilities across multiple trials, we perform $N$ independent rerankings of the candidate list $\textbf{I}_{\text{can}}$. The resulting sequences are then aggregated using a simplified variant of Reciprocal Rank Fusion (Chapelle et al. 2009). For each item $i_{k}\in\textbf{I}_{\text{can}}$, we compute an aggregated score as follows:
$g(i_{k})=\sum_{n=1}^{N}\frac{1}{\varsigma^{(n)}(i_{k})+1}$ (3)
where $\varsigma^{(n)}(i_{k})$ denotes the position of item $i_{k}$ in the $n$-th reranked sequence. We select the top $p$ items with the highest scores to construct new artificial interactions for the user $u$. We repeat the process for all targeted users, leading to $Q_{\alpha}\times|\mathcal{V}|\times p$ newly augmented edges. After data augmentation, we obtain an enriched graph $\mathcal{G}_{\text{aug}}=(\mathcal{V},\mathcal{E}^{+})$, where the original edge set $\mathcal{E}$ is augmented with newly generated interactions, thereby expanding $\mathcal{E}^{+}=\mathcal{E}\cup\mathcal{E}_{\text{new}}$ while preserving the bipartite structure between users $\mathcal{U}$ and items $\mathcal{I}$.
Theorem 1 (Concentration for Score Aggregation over Random Permutations).
Let =$\mathcal{I}=\{i_{1},i_{2},\ldots,i_{K}\}$ be a set of $K$ items, and let $\mathcal{T}_{K}$ denote all permutations of $\{1,2,\ldots,K\}$. Let $P$ be any probability distribution over $\mathcal{T}_{K}$, and let $\varsigma^{(1)},\ldots,\varsigma^{(N)}$ be independent random permutations sampled from $P$. Let $g:\mathbb{R}\to\mathbb{R}$ be a bounded, strictly decreasing function. Suppose $A,B\in\mathbb{R}$ are such that $A\leq g(x)\leq B$ for all $x$. For each item $i\in\mathcal{I}$, define the aggregated score
$S(i)=\sum_{n=1}^{N}g\big{(}\varsigma^{(n)}(i)\big{)}.$
Fix distinct items $i_{j},i_{k}\in\mathcal{I}$, and let
$\mu=\mathbb{E}_{\varsigma\sim P}\left[g\big{(}\varsigma(i_{j})\big{)}-g\big{(}\varsigma(i_{k})\big{)}\right].$
If $\mu<0$, then
$\Pr\big{(}S(i_{j})>S(i_{k})\big{)}\leq\exp\left(-\frac{N\mu^{2}}{2(B-A)^{2}}\right).$
This theorem is sufficent for $g(x)$ in Equation 3 and $\forall i_{j},i_{k}\in\mathcal{I}$, let $\mu=\mathbb{E}\left[\frac{1}{\varsigma(i_{j})+1}-\frac{1}{\varsigma(i_{k})+1}\right].$ If $\mu<0$, then
$\Pr\big{(}S(i_{j})>S(i_{k})\big{)}\leq\exp\left(-\frac{N\mu^{2}}{2\left(1-\frac{1}{K+1}\right)^{2}}\right).$
Theorem 1 illustrates that with a sufficiently large number of LLM reranking steps $N$, the ensemble output converges toward the LLM’s optimal parametric knowledge for the reranking task, thereby yielding high-confidence augmented data. We provided detailed proof of Theorem 1 in Appendix LABEL:theorem.
Method Virtual Few-shot Textual Consistent
Interaction Prompting Embed. Aug. Data
VoteGCL ×
LLMRec × ×
(Wei et al. 2024)
RLMRec × × ×
(Ren et al. 2024)
KAR × × ×
(Xi et al. 2024)
Table 1: Comparison of LLM augmented data methods
Unlike other LLM-augmented data approaches summarized in Table 1, VoteGCL does not incur the cost of generating textual embeddings for user/item profiling or adapting high-dimensional textual representations into recommendation models. Instead, VoteGCL focuses on generating high-quality interaction data and effectively leveraging it through a Graph CL paradigm, as detailed in Section 4.2. The key advantages of our LLM-augmented data are twofold: (1) a majority-vote reranking mechanism, which ensures more consistent and high-confidence user-item interactions, with detail theoretical justification provided in Appendix LABEL:theorem; and (2) a few-shot prompting strategy that incorporates collaborative signals from similar users, enabling better adaptation to user-specific preferences in the data augmentation task.
4.2 Graph Contrastive Learning Paradigm
We continue to employ LightGCN as our graph encoder due to its efficiency and effectiveness, though our framework is compatible with other graph backbones (e.g., NGCF (Wang et al. 2019), XSimGCL (Yu et al. 2023)). Inspired by SimCLR (Chen et al. 2020), we employ a CL paradigm between two representation views derived from the original and augmented data. This approach mitigates distributional shift caused by potential noise introduced during graph augmentation. We share a common embedding matrix $\mathbf{E}^{(0)}$ to encode both the original graph $\mathcal{G}$ and its augmented version $\mathcal{G}_{\text{aug}}$. During each training epoch, node representations are generated as follows:
$\mathbf{E}^{(l)}=H(\mathbf{E}^{(l-1)},\mathcal{G}_{\text{aug}}),\quad\mathbf{E}_{\text{org}}^{(l)}=H(\mathbf{E}^{(l-1)},\mathcal{G})$
These embeddings serve as two distinct views of representation learning for contrastive supervision. The contrastive loss is then applied by treating embeddings of the same node across different views as positive pairs, and all others as negative:
$\mathcal{L}_{\text{cl}}=\sum_{i\in\mathcal{B}}-\log\frac{\exp\left(\mathbf{z}_{i}^{\top}\mathbf{z}_{i,\text{org}}/\tau\right)}{\sum_{j\in\mathcal{B}}\exp\left(\mathbf{z}_{i}^{\top}\mathbf{z}_{j,\text{org}}/\tau\right)},$ (4)
Here, $i,j$ represent user (item) nodes within a sampled mini-batch $\mathcal{B}$, and $\mathbf{z},\mathbf{z}_{\text{org}}$ are the $L_{2}$-normalized $d$-dimensional representations obtained from the augmented graph $\mathcal{G}_{\text{aug}}$ and original graph $\mathcal{G}$. Note that, $\mathbf{z}_{i}=\frac{\mathbf{e}_{i}}{\left\|\mathbf{e}_{i}\right\|_{2}}$. By replacing $\tilde{A}$ with the adjacency matrices of graph $\mathcal{G}_{\text{aug}}$, $\mathbf{z}$ can be learned via Equation 1. The temperature parameter $\tau>0$ (e.g., $\tau=0.2$) controls the sharpness of the similarity distribution. The CL loss can also be expressed as:
$\mathcal{L}_{cl}=\sum_{i\in\mathcal{B}}\left(-\mathbf{z}_{i}^{\top}\mathbf{z}_{i,\text{org}}/\tau+\log(\sum_{j\in\mathcal{B}/\{i\}}\exp(\mathbf{z}_{i}^{\top}\mathbf{z}_{j,\text{org}}/\tau))\right)$ (5)
Thus, $\mathcal{L}_{\text{cl}}$ enhances uniformity in the embedding space through its first component and promotes diversity through the second component, thereby improving model accuracy and mitigating popularity bias (Zimmermann et al. 2021; Wang and Isola 2020). The final node embeddings used for recommendation are obtained from the last propagation layer of the augmented graph. Specifically, we adopt $\mathbf{E}^{(L)}$, where $L$ corresponds to the total number of propagation layers, as the learned representation from the augmented graph $\mathcal{G}_{\text{aug}}$. This embedding captures the high-order structural information introduced by the augmentation and is subsequently employed to compute the BPR loss for optimizing recommendation performance. Finally, the overall training objective combines the recommendation loss with the CL loss:
$\mathcal{L}_{\text{main}}=\mathcal{L}_{\text{BPR}}+\lambda\mathcal{L}_{\text{cl}}$ (6)
where $\lambda\in(0,1)$ is a hyperparameter to control weight of CL loss. The main advantages of this CL paradigm are that it leverages LLMs purely for knowledge-guided augmentation without embedding integration, thereby avoiding high-dimensional representations and complex fusion with collaborative signals. Additionally, VoteGCL shares the same learnable embedding matrix for both the original and augmented graphs, enabling more efficient training with reasonable model complexity. A detailed analysis of model complexity is provided in Appendix LABEL:complex.
5 Experiments
5.1 Experimental Setup
Dataset. We evaluate VoteGCL on four real-world datasets: Amazon (Scientific and Book) (Hou et al. 2024a), MovieLens [^1] https://grouplens.org/datasets/movielens/, Netflix (Wei et al. 2024), and Yelp2018 (Wang et al. 2019). The dataset statistics after preprocessing are summarized in Table 2. By default, we select the top-1 result from majority-vote LLM reranking as virtual edges for data augmentation.
Datasets Users Items Inters. Density Aug.
Amazon Scientific 10,995 5,314 76,091 $1.3\times 10^{-3}$ 10,995
Amazon Book 11,000 9,332 212,484 $2.1\times 10^{-4}$ 11,000
Movielen-100K 610 9,701 100,836 $1.7\times 10^{-2}$ 610
Movielen-1M 6000 25,317 914,108 $6.2\times 10^{-3}$ 6000
Yelp2018 11,091 86,646 544,672 $5.3\times 10^{-4}$ 11,091
Netflix 13,187 17,336 68,933 $3.2\times 10^{-4}$ 13,187
Table 2: Dataset statistics
Experimental Setup. We use NDCG@K and Recall@K to evaluate the alignment between recommended items and user preferences, reflecting ranking quality and retrieval accuracy. APLT@K (Abdollahpouri, Burke, and Mobasher 2019) measures the average percentage of long-tail items among the top-K recommendations, serving as an indicator of popularity bias. Following the leave-one-out strategy, we use the latest interaction for testing, the second latest for validation, and the rest for training. Metrics are computed over all items, and average scores are reported. All models are trained for 100 epochs with 256-dimensional embeddings and a learning rate of $1\text{e}^{-3}$; baseline hyperparameters follow their original settings. LightGCN is used as the retrieval model to generate item candidates for LLM reranking. For augmentation, we use the same LLM as prior work gpt-3.5-turbo (Achiam et al. 2023). All experiments are conducted on a single NVIDIA A5000 GPU.
Baselines. We compare VoteGCL against three categories of baseline models: (1) base graph-based recommendation systems, including LightGCN (He et al. 2020), NGCF (Wang et al. 2019), and MF-BPR (Koren, Bell, and Volinsky 2009); (2) traditional augmented graph recommendation methods, such as SGL (Zhou et al. 2023), GraphDA (Fan et al. 2023b), GraphAug (Zhang et al. 2024), and XSimGCL (Yu et al. 2022); and (3) LLM-enhanced recommendation systems, including KAR (Xi et al. 2024), RLMRec (Ren et al. 2024), and LLMRec (Wei et al. 2024). Detailed descriptions of the baselines are provided in Appendix LABEL:experiment.
5.2 Main results
Dataset Metric MF-BPR LightGCN NGCF GraphAug GraphDA SGL XSimGCL
NGCF+
VoteGCL
XSimGCL+
VoteGCL
LightGCN+
VoteGCL
Improv.
Movielen
100K
R@10 0.0750 0.0766 0.0681 0.0689 0.0528 0.0818 0.0852 $0.0691^{0.3\%}$ $0.0852^{0.0\%}$ $\textbf{0.0886}^{15.7\%}$ 3.99%
N@10 0.0335 0.0334 0.0308 0.0334 0.0251 0.0407 0.0398 $0.0372^{8.1\%}$ $0.0411^{3.3\%}$ $\textbf{0.0494}^{47.9\%}$ 21.37%
A@10 0.0042 0.0097 0.0019 0.0016 0.0463 0.0227 0.0583 $0.0166^{733\%}$ $\textbf{0.1852}^{217\%}$ $0.0617^{536\%}$ 217.66%
Movielen
1M
R@10 0.0712 0.0817 0.0628 0.0643 0.0799 0.0812 0.0857 $0.0659^{4.94\%}$ $0.0901^{5.13\%}$ $\textbf{0.0928}^{13.6\%}$ 8.28%
N@10 0.0365 0.0403 0.0311 0.0268 0.0385 0.0402 0.0416 $0.0375^{20.6\%}$ $0.0428^{2.88\%}$ $\textbf{0.0462}^{14.6\%}$ 11.06%
A@10 0.0015 0.0027 0.0011 0.0056 0.0104 0.0028 0.0098 $0.0061^{454\%}$ $0.0123^{25.5\%}$ $\textbf{0.0145}^{437\%}$ 39.42%
Yelp2018 R@10 0.0304 0.0425 0.0375 0.0352 0.0422 0.0446 0.0435 $0.0401^{6.9\%}$ $0.0442^{1.61\%}$ $\textbf{0.0487}^{14.6\%}$ 8.79%
N@10 0.0147 0.0214 0.0229 0.0125 0.0230 0.0223 0.0221 $0.0201^{0.0\%}$ $0.0234^{10.9\%}$ $\textbf{0.0247}^{15.4\%}$ 7.86%
A@10 0.0075 0.0038 0.0014 0.0012 0.0032 0.0096 0.0068 $0.0018^{28.6\%}$ $0.0210^{208\%}$ $\textbf{0.0349}^{818\%}$ 263.54%
Amazon
Scientific
R@10 0.1173 0.1251 0.0941 0.0986 0.1204 0.1297 0.1291 $0.1317^{39.9\%}$ $0.1335^{3.4\%}$ $\textbf{0.1385}^{10.7\%}$ 6.78%
N@10 0.0850 0.0853 0.0648 0.0646 0.0758 0.0911 0.0896 $0.1008^{48.0\%}$ $0.0936^{4.5\%}$ $\textbf{0.0993}^{16.4\%}$ 9.01%
A@10 0.2563 0.2015 0.1839 0.0413 0.2034 0.2504 0.1779 $0.2542^{38.2\%}$ $0.1975^{11.0\%}$ $\textbf{0.2795}^{38.7\%}$ 9.05%
Table 3: Performance Comparison of Graph Recommendation Methods on Multiple Datasets Before and After Data Augmentation. Each value in augmented data is shown as $\text{value}^{\text{improvement}}$, where the superscript indicates the percentage improvement over the baseline. Underlined values highlight the best baseline performance, while bold values highlight the best performance after data augmentation. ”Improv” refers to the absolute improvement between these two.
Figure 5: Performance of LLM-enhanced graph recommenders on three datasets under @10 and @20 settings, using dataset-specific models and leveraging available LLM-augmented data for fair comparison.
Traditional graph recommendation
[…]
LLM-enhanced graph recommendation system
As shown in Figure 5, LightGCN + VoteGCL consistently outperforms all other methods across three benchmark datasets (Netflix, Amazon Book, and Yelp2018) under both @10 and @20 evaluation settings, achieving superior performance in terms of both accuracy and diversity. This includes outperforming traditional baselines (LightGCN, SGL) as well as LLM-enhanced models. While LLMRec, RLMRec, and KAR demonstrate notable improvements in ranking metrics such as NDCG particularly when textual embeddings from LLMs are incorporated into the recommendation backbone, they also introduce a higher degree of popularity bias. This drawback is reflected in their lower APLT scores compared to SGL, which applies conventional data augmentation without relying on LLM-generated semantics. In contrast, VoteGCL alleviates this issue by generating high-confidence synthetic interactions through LLM prompting and integrating them effectively via graph contrastive learning, resulting in a more balanced trade-off between recommendation accuracy and fairness.
5.3 Ablation Studies
Number of Virtual Edges for Augmentation
[…]
Figure 6: Comparison of reranking performance between zero-shot and few-shot prompting across different numbers of majority-voting rounds. Results are averaged over 30 runs on the Amazon Scientific and MovieLens-100K datasets.
Study of VoteGCL
Figure 7: Performance of VoteGCL under different contrastive weight
Dataset Metric Quantile of users to be augmented
[…]
MovieLens-100K N@10 0.0494 0.0482 0.0501 0.0503
A@10 0.0617 0.0562 0.0574 0.0596
Amazon Scientific N@10 0.0993 0.0983 0.0989 0.0991
A@10 0.2795 0.2939 0.2893 0.2792
Netflix N@10 0.0351 0.0326 0.329 0.0342
A@10 0.0892 0.1089 0.0995 0.0989
Table 4: Performance under different ratios of users selected for augmentation
Methods MovieLens-100K Netflix Amazon Book
N@10 A@10 N@10 A@10 N@10 A@10
VoteGCL 0.0494 0.0617 0.0351 0.0892 0.0781 0.3654
LightGCN 0.0344 0.0097 0.0221 0.0358 0.0486 0.2390
w/o GCL 0.0362 0.0102 0.0235 0.0341 0.0488 0.2511
w/o LLM 0.0201 0.0068 0.0115 0.0126 0.0274 0.1243
Table 5: Ablation Study on VoteGCL Components
Table 5 presents an ablation study evaluating the impact of key components in VoteGCL across three datasets. The full VoteGCL model consistently outperforms all baselines in both NDCG@10 and APLT@10. When the graph CL module is removed (”w/o GCL”), performance drops across all datasets, though it still outperforms LightGCN in certain metrics. This suggests that while the LLM-based augmentation is beneficial, without GCL the model becomes more sensitive to noise from the synthetic data. Additionally, removing the LLM reranking with majority voting (”w/o LLM”) and selecting top-1 candidates without using LLM to vote leads to a significant performance collapse. This highlights the importance of majority voting in stabilizing the augmentation and leveraging textual descriptions modalities, making both GCL and LLM-based reranking essential for the effectiveness of VoteGCL.
Zero-shot vs Few-shot Prompting LLM with Different Number of Majority-voting
Figure 6 illustrates the impact of prompting strategy and the number of reranking rounds on the NDCG@10 performance of LLM-based item reranking. Across both datasets, few-shot prompting consistently outperforms zero-shot prompting, highlighting its effectiveness in capturing user preferences through collaborative signals. As the number of majority-voting rounds increases, performance improves and becomes more stable, particularly from 8 rounds onward, for both prompting strategies. Notably, the few-shot setting yields higher mean scores across 30 trials, indicating greater reliability and robustness in reranking outcomes. These findings suggest that effectively leveraging the LLM’s parametric knowledge via few-shot prompting leads to higher-quality augmented data and ultimately enhances the performance of VoteGCL. Notably, the performance at 6 voting rounds is already competitive, making it a practical and cost-effective choice in resource-constrained scenarios. Based on these observations, we recommend combining few-shot prompting with a sufficiently large number of majority-voting rounds (e.g., 6$\geq$ 6≥ 6) to fully realize the benefits of LLM-based reranking in data augmentation.
Controling Hyperparam $\lambda$ Contrastive Weight
As illustrated in Figure 7, the contrastive loss weight $\lambda\in(0,1)$ plays a critical role in balancing recommendation accuracy and popularity bias mitigation. When $\lambda$ increases, the APLT@10 metric improves significantly, indicating greater inclusion of long-tail items and enhanced diversity in recommendations. This effect is consistent across both MovieLens-100K and Netflix datasets. However, this gain in diversity comes with a trade-off NDCG@10 gradually declines as $\lambda$ grows, reflecting a reduction in ranking precision. The results suggest that overly high values of $\lambda$ (e.g., 0.2) may overly prioritize long-tail exploration at the expense of accuracy. Based on these observations, we recommend tuning $\lambda$ within the range of 0.01 to 0.1 to strike an effective balance between accuracy and fairness, depending on the characteristics of the target dataset.
Additional Experimental Results
Additional ablation studies are provided in Appendix LABEL:experiment. These include evaluations of different retrieval models for generating candidate items for LLM reranking, assessments of alternative LLMs for data augmentation (e.g., Qwen2.5 (Yang et al. 2024)), and an analysis of how varying the number of top-K items in the majority-voting process affects performance. Furthermore, we compare the cost of generating augmented data using VoteGCL against other LLM-based methods, demonstrating that our approach offers a cost-efficient solution.
6 Conclusion
We propose VoteGCL, a model-agnostic data augmentation framework that tackles inconsistency and distributional mismatch in LLM-enhanced recommendation. VoteGCL uses LLM-based reranking with majority voting to generate reliable synthetic interactions and applies contrastive learning to align original and augmented graphs without altering model architecture. Experiments across benchmarks show VoteGCL improves recommendation accuracy and reduces popularity bias, outperforming traditional and LLM-based data augmentation methods for graph-based recommender.